The perimeter of rectangle ABCD is 40 ft. If BC is 3 times as long as CD, what is the length of AD? (BC and CD are suppose to have lines on top of them, and ABCD are points on each corner of a rectangle).
Answer by sachi(548)

let CD =x then BC = 3x
so perimeter= =2(x+3x)=8x=40ft
so x=40/8=5=Ab=CD
& AD=BC=3*5=15 ft ans